# PPR: Enhancing Dodging Attacks while Maintaining Impersonation Attacks on Face Recognition Systems

### v2

@@ 4.1 Experimental Setting
For calculating the attack success rate in impersonation and dodging attacks, we choose the thresholds based on FAR@0.001 on the entire LFW dataset.
Attack Setting. Without any particular emphasis, we set the maximum allowable perturbation magnitude, denoted as $\epsilon$, to 10. This value is based on the $L_{\infty}$ norm bound, considering pixel values within the range of $[0,255]$. Furthermore, we specify the maximum number of iterative steps as 200. Moreover, both $\beta$ and $\gamma$ are set to 1.0.
Evaluation Metrics. We employ the attack success rate (ASR) to assess the performance of various attacks. ASR represents the ratio of successfully attacked adversarial face examples.
Baseline methods Our proposed attack is a restricted attack method that aims to enhance the robustness of FR models. It is not rational to compare our proposed method with the unrestricted attacks which do not limit the magnitude of the adversarial perturbations. Therefore, we choose restricted attacks on FR that aim to enhance the robustness of FR models(Zhong & Deng, 2021)(Zhou et al., 2023b) and transfer attacks(Xie et al., 2019)(Long et al., 2022)(Wang et al., 2023b)(Wang et al., 2023a) as our baseline.
Table 2: Comparisons of dodging ASR with multi-task attacks on the LFW dataset. The models in the first row are the victim models.
[…]
@@
Lagrangian 3.9 26.5 100.0 26.0
Lagrangian + ours 7.3 36.4 100.0 26.6
DA 11.0 35.6 99.1 37.4
DA + ours 17.5 44.9 99.4 37.8
4.2 The Trade-off Between the Impersonation Attacks and Dodging Attacks
Owing to the inherent conflict during the optimization process of impersonation and dodging losses, there exists a trade-off between impersonation and dodging performance. We craft adversarial face examples using Lagrangian attack and our proposed PPR on LFW based on DI. The average black-box results are demonstrated in Figure 2.
Figure 2: The trade-off between the performance of impersonation attack and dodging attack of adversarial examples.
Figure 2 illustrates that the inclusion of our proposed method into the Lagrangian attack results in reduced the trade-off. It is not hard to see the reason. The pruning operation of our proposed PPR serves to sparsify the adversarial perturbations while preserving the impersonation performance. On the other hand, the restoration operation tends to introduce adversarial perturbations in the pruned areas, specifically favoring dodging attacks. These operations effectively enhance the dodging attack performance while maintaining the impersonation attack performance, ultimately mitigating the trade-off.
4.3 Comparison Study
We compare our proposed attack method with the baseline attacks on multiple FR models and datasets. For our proposed method, we choose the Lagrangian attack as the attack for both the stages of pre-training and restoration. The results are shown in Table 1. The numbers before the slash represent the results of the baseline attack, which refers to the attack method mentioned in the seconds column that solely utilizes the traditional impersonation attacks to create the adversarial examples. On the other hand, the numbers after the slash correspond to the results obtained by combining the baseline attack with our proposed PPR.
Table 1 shows that after being added our proposed attack method, the dodging ASR of the adversarial attacks improve a lot. Note that the average black-box impersonation ASRs of the baseline attacks in Table 1 increase after incorporating our proposed attack method. This demonstrate the effectiveness of our proposed method in improving the dodging attack performance while maintain the impersonation attack performance.
[…]
In practical application scenarios, victims can employ adversarial robust models to defend against adversarial attacks. Consequently, it becomes crucial to evaluate the performance of adversarial attacks on these robust models. In this study, we generate adversarial examples on the LFW dataset using MF as the surrogate model and assess the performance of various attacks on the adversarial robust models. The results are presented in Table 3.
Table 3 illustrates that the inclusion of our proposed method leads to improvements in both dodging and impersonation performance. These results serve as evidence of the effectiveness of our proposed method on adversarial robust models.
Table 3: Comparisons of dodging ASR on LFW with adversarial robust models as victim models.
[…]
DI 2.3 7.8 0.5 19.8
DI + ours 7.2 17.1 4.7 20.3
VMI 0.8 4.5 0.3 15.8
VMI + ours 3.3 8.9 3.0 16.2
SSA 0.6 2.9 0.1 14.7
SSA + ours 3.9 9.7 3.0 14.8
DFANet 0.2 0.8 0 8.1
DFANet + ours 0.3 1.3 0.1 8.3
SIA 1.4 6 0.2 17.5
SIA + ours 5.6 15.5 5.6 17.6
BSR 0.4 1.9 0 6.4
BSR + ours 0.7 3.5 0.4 6.6
BPFA 0.3 1.8 0.1 13.1
BPFA + ours 4.1 8.4 4.2 13.4
4.4 Ablation Study
To delve deeper into the properties of our proposed attack method, we conducted an ablation experiment using DI as the Baseline attack, with MF serving as the surrogate model on the LFW dataset. To confirm the effectiveness of our pruning method, we employed the Random Zeroing (RZ) method, which randomly sets adversarial perturbations to zero. We applied this method and our pruning method to free up 20% of the adversarial perturbations. For Fine-tuning, we employed the Lagrangian attack as the method to further optimize the adversarial examples crafted by the Lagrangian attack. The dodging attack ASR and average black-box impersonation ASR results are shown in Table 4.
Table 4: Comparisons of ASR results of dodging attack and impersonation attacks on the LFW dataset. * denote the method that only use the impersonation loss to craft the adversarial face examples.
[…]
@@ 4.4 Ablation Study
Baseline 2.2 18.2 99.2 25.6
Lagrangian 3.9 26.5 100.0 26.0
Fine-tuning 3.4 23.2 99.9 25.8
RZ 0.6 4.4 94.8 15.1
Pruning 3.4 24.8 100.0 25.4
PPR 6.8 35.5 100.0 25.9
Table 4 demonstrates that our proposed pruning method for adversarial examples achieves a significantly smaller decrease in ASR than RZ after pruning 20% of adversarial perturbations, indicating the effectiveness of our proposed approach. After pruning and restoration, adversarial perturbations using our proposed Pruning method, both impersonation and dodging ASR of the crafted adversarial face examples are recovered and higher than the Baseline attack method. These results demonstrate the effectiveness of our proposed PPR in improving the dodging performance of adversarial attacks on FR without compromising the impersonation attack performance.
Figure 3: The dodging ASR in different sparsity ratio.
4.5 Hyperparameter Sensitivity Study
The sparsity ratio quantifies the proportion of adversarial perturbations that are allowed to be discarded during the pruning stage. This ratio greatly impacts the performance of our proposed attack method. Hence, we conducted a sensitivity study on the sparsity ratio to analyze its effect on performance of the algorithm. We selected the DI-based Lagrangian attack method as the Baseline for pre-training and performed a hyperparameter sensitivity study on LFW with FaceNet as the surrogate model. We adjusted the value of $\tilde{\lambda}$ to ensure that the average black-box impersonation ASR results were within a 0.8% absolute difference compared to the corresponding Baseline. The dodging ASR results are illustrated in Figure 3.
Figure 3 illustrates that the dodging ASR of our proposed method initially increases and then decreases as the sparsity ratio increases. It is not hard to see the reason; When the sparsity ratio increases, a greater number of adversarial perturbations are pruned, creating more empty regions for the adversarial perturbations that favor dodging attacks in the restoration stage. If the sparsity ratio is set to a too high value, an excessive number of adversarial perturbations are allowed to be freed up, resulting in a degradation of performance for the pre-trained adversarial examples. Consequently, the performance of adversarial face examples will decrease.
5 Conclusion
In this paper, we delve into the issue of multi-identity samples present among adversarial face examples. Our research reveals that traditional impersonation attacks often involve multi-identity samples and the success of an traditional impersonation attack may not necessarily imply success in dodging attacks in the black-box setting. In order to improve dodging performance without compromising impersonation performance, we proposed a novel attack known as PPR. PPR involves pruning the adversarial examples to release some of the adversarial perturbations while tending to maintain the performance of the adversarial examples. We then restore the adversarial perturbation in the pruned area, which favors dodging attack, to enhance the performance of dodging attack without harming the performance of the impersonation attack. Our extensive experiments demonstrate the effectiveness of our proposed attack method.
6 Potential Broader Impact
This paper showcases research dedicated to advancing the field of Machine Learning. The proposed attack method could jeopardize the security of FR systems. Our aim is to raise awareness through our proposed attack method and reinforce the robustness of FR systems.
Appendix A Multi-identity Samples among the Benign Face Images
Multi-identity samples are intriguing samples that can be classifies as multiple classes in FR. In this subsection, we will explore the existence of multi-identity samples among the benign face images (i.e. the face images without being added adversarial perturbations).
We randomly select negative face pairs from the entire LFW dataset. Subsequently, we use MF as our FR model to extract the embeddings of the face images in each face pair and calculate the cosine similarity between the two images. If the cosine similarity surpasses the threshold, both images in the pair are classified as belonging to multiple identities, indicating that they are multi-identity samples. Our findings demonstrate the presence of multi-identity samples among the benign face images, as illustrated in Figure 4.
Figure 4 illustrates that the multi-identity samples among the benign face images closely resemble the appearances of the identities they are classified into, suggesting they are hard samples(Zhou et al., 2023b)(Huang et al., 2020).
To analyze the cause of this phenomenon, we need to consider the properties of both the multi-identity samples and the FR model. Commonly-used FR models are well-trained and capable of correctly classifying the majority of benign face images. However, there are some benign face images that the FR model fails to classify accurately, and these samples are referred to as hard samples. Multi-identity samples are a specific type of hard sample known as hard negative samples. Typically, hard negative samples exhibit a similar appearance(Zeng et al., 2020), indicating that the multi-identity samples among the benign face images share a resemblance.
Figure 4: The Illustration of the multi-identity samples. The images in the first row and third row are the images of different identities. The image in the second row are the images of the multi-identities samples that will be recognized as the identities both top and below it by the FR model.
Appendix B Universality of Multi-identity Samples among Adversarial Face Examples
The traditional impersonation attacks craft adversarial face example only use the impersonation loss (e.g. Equation 5). In this subsection, we will investigate the ratio of multi-identity samples and evaluate the effectiveness of traditional impersonation attacks in terms of dodging attacks. We utilize the multi-identity sample ratio (MSR) to gauge the proportion of multi-identity samples that can be recognized as both the attacker and victim identities in the adversarial face examples capable of executing successful impersonation attacks. MSR be calculate as:
${\rm MSR}=\frac{|\mathcal{S}_{1}\cap\mathcal{S}_{2}|}{|\mathcal{S}_{2}|}$ (18)
where $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ are the set of the adversarial examples that are identified as the attacker and victim identities, respectively. Therefore, the $\mathcal{S}_{2}$ is the set of the adversarial face examples that can perform successful impersonation attack. Equation 3 can be rewritten as:
${\rm ASR}^{i}=\frac{|\mathcal{S}_{2}|}{N_{p}}$ (19)
We evaluate the MSR, impersonation attack ASR and dodging attack ASR using MF as the surrogate model on the LFW dataset. The results are shown in Table 5.
@@ Appendix B Universality of Multi-identity Samples among Adversarial Face Examples
Table 5: The results (%) on the LFW dataset. The models on the first row are the victim models.
@@
[…]
Table 5 demonstrate that most of the crafted adversarial face examples are multi-identity samples when the victim model are the black-box model. This indicates that most adversarial face examples generated through traditional impersonation attacks are unable to attain a successful dodging attack in the black-box setting. Nevertheless, in the white-box setting, the majority of adversarial face examples capable of executing successful impersonation attacks also demonstrate success in dodging attacks.
To analyze the reason, we need to consider the metric used to determine whether two face images belong to the same identity in FR. FR is an open-set task, so we utilize the distance in the embedding space to make this determination. Figure 5 illustrates two decision boundaries for each FR model, one for attacker identity and one for victim identity.
In the white-box setting, if we generate adversarial examples using Equation 5, these examples can enter a space in which they are identified as the victim identity but not as the attacker identity. However, the decision boundary of the black-box model differs from that of the surrogate model. In the black-box setting, most adversarial examples are found between the decision boundaries of the black-box model, resulting in the majority of adversarial examples crafted using Equation 5 being multi-identity samples. This demonstrates that the adversarial face examples are positioned near the decision boundary, aligning with findings from prior research of adversarial attacks on image classification(Cao & Gong, 2017).
Figure 5: Illustration of the decision boundaries of FR models. The star in the left and right are the initial image of the adversarial face examples (i.e. the attacker image) and the final adversarial example, respectively. The solid and dashed lines are the decision boundaries for the surrogate and black-box models, respectively. For both the surrogate and black-box models, the lines in the left and right are the decision boundary of the victim and attacker, respectively.

### v3

@@ Appendix C The proof of the supremum of the absolute variance
Let us revisit an early hypothesis posited by Goodfellow et al. [14], suggesting that the linear nature of modern DNNs, resembling linear models trained on the same dataset, is the underlying cause of adversarial examples and their surprising transferability [15]. Based on the hypothesis, the FR model $\mathcal{F}$ can be represented in the following linear form:
$\mathcal{F}\left(x\right)=\omega x+b$ (16)
where $\omega$ is the weight vector, and $b$ is the bias term.
Hence, the absolute variance of the output of the function $\mathcal{F}$ can be articulated as:
$\begin{split}\rho&=\|\mathcal{F}\left(x+\delta\right)-\mathcal{F}\left(x\right)\|\\ &=\|\omega\left(x+\delta\right)+b-\left(\omega x+b\right)\|\\ &=\|\omega\delta\|\leq\|\omega\|\|\delta\|=\|\omega\|\mathcal{I}\\ \end{split}$ (17)
The supremum of the absolute variance is given by $\lVert\omega\rVert\mathcal{I}$. Within the context of gradient-based adversarial attacks, the parameters of the neural networks are typically held constant. Consequently, the vector $\omega$ can be treated as a constant. From this, we can deduce that the supremum of the absolute variance in the output of $\mathcal{F}$ is directly proportional to the priority measure $\mathcal{I}$. It follows that adversarial perturbations of greater magnitude are more prioritized, as they exert a more pronounced influence on the output of the face recognition model $\mathcal{F}$.
Appendix D The Calculation Method of the Attack Success Rate
Due to the different objectives of the impersonation and dodging attacks, the calculation methods for the ASR also vary between the two attack types. When it comes to impersonation attacks on FR, the ASR can be computed as:
where the value of $t^{d}$ represents the threshold for dodging attacks.
Appendix E The Calculation Method of the Multi-identity Sample Ratio
Multi-identity Sample Ratio (MSR) is the proportion of multi-identity samples that can be recognized as both the attacker and victim identities in the adversarial face examples capable of executing successful impersonation attacks. The detailed calculation method of MSR be expressed as:
${\rm MSR}=\frac{|\mathcal{S}_{1}\cap\mathcal{S}_{2}|}{|\mathcal{S}_{2}|}$ (20)
where $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ are the set of the adversarial examples that are identified as the attacker and victim identities, respectively. Therefore, the $\mathcal{S}_{2}$ is the set of adversarial face examples that can perform a successful impersonation attack.
Table 4: Comparisons of ASR results (%) of dodging attack and impersonation attacks on the LFW dataset using the negative cosine similarity loss. The models in the second row are the victim models.
ASRd ASRi
@@ Appendix E The Calculation Method of the Multi-identity Sample Ratio
Baseline 2.2 16.1 99.4 25.4
Baseline + Lagrangian 4.0 26.5 100.0 26.0
Baseline + ours 8.1 39.7 100.0 26.2
Appendix F Results on Negative Cosine Similarity Loss
Apart from the normalized $L_{2}$ distance loss (i.e. Eq. 1 and Eq. 2), the negative cosine similarity loss is also a commonly used loss function in adversarial attacks on FR systems. In this section, we will assess the effectiveness of our proposed method regarding the negative cosine similarity loss.
For impersonation attacks, the negative cosine similarity loss can be expressed as:
$\mathcal{L}^{i}_{c}=1-\cos\left(\mathcal{F}\left(x^{adv}\right),\mathcal{F}\left(x^{t}\right)\right)$ (21)
where $\cos\left(e,e^{\prime}\right)$ is the function to calculate the cosine similarity between $e$ and $e^{\prime}$.
Similarly, the negative cosine similarity loss for dodging attacks can be formulated as follows:
$\mathcal{L}^{d}_{c}=1+\cos\left(\mathcal{F}\left(x^{adv}\right),\mathcal{F}\left(x^{s}\right)\right)$ (22)
To evaluate the effectiveness of our proposed method on the negative cosine similarity loss, we use DI as the Baseline to craft adversarial face examples using the negative cosine similarity loss $\mathcal{L}^{i}_{c}$ and $\mathcal{L}^{d}_{c}$ as $\mathcal{L}^{i}$ and $\mathcal{L}^{d}$, respectively. The performance is demonstrated in Table 4.
Table 4 demonstrates that the performance of both impersonation and dodging improves after integrating our proposed attack method into the Baseline attack. This further highlights the effectiveness of our method regarding the negative cosine similarity loss.
Table 5: The results (%) on the LFW dataset using MF as the surrogate model. The models on the first row are the victim models.
ASRd ASRi
@@ Appendix F Results on Negative Cosine Similarity Loss
SA 99.4 45.9 2.1 44.5
SA + Lagrangian 100.0 49.2 3.4 44.8
SA + ours 100.0 54.9 4.7 46.4
Appendix G Results on Sibling-Attack
Sibling-Attack [29] (abbreviated as SA) is a cutting-edge adversarial attack on FR that demonstrates considerable effectiveness in enhancing the transferability of adversarial attacks on FR. SA employs a surrogate model with a two-branch multi-task neural network, which differs from the surrogate models discussed in Section 4.2. Consequently, we exclusively consider SA for comparison with our method.
We aligned the attack setting with the official SA implementation and integrated our proposed Adv-Pruning with SA. The FFHQ dataset, provided by the official SA page [^1] https://github.com/Tencent/TFace/tree/Adv-Attack-Defense/security/tasks/Adv-Attack-Defense/Sibling-Attack, is chosen as the dataset for generating adversarial face examples. The results are depicted in Table 5.
Table 5 demonstrates that the incorporation of our proposed method results in enhancements in both dodging and impersonation performance. This further demonstrates the effectiveness of our proposed attack method in enhancing the dodging attack performance without compromising the impersonation attack performance.
@@ Appendix G Results on Sibling-Attack
Table 6: The results (%) on the LFW dataset using MF as the surrogate model. The models on the first row are the victim models.
@@
ASRi 15.8 17.8 100
ASRd 0.9 4.6 97.7
Appendix H Results on MSR of Previous Impersonation Attacks
Following the same experimental setting of Section 4.4, we conduct experiments on various previous impersonation attacks. The results are shown in Table 6.
As indicated by Table 6, the majority of the crafted adversarial face examples are multi-identity samples in the black-box setting. This observation suggests that the majority of adversarial face examples, which are generated through the previous impersonation attacks, fail to achieve a successful dodging attack in the black-box setting.